Global insurance has found that 20% (1 in 5) of all insurance policies are surrendered (cashed in) before their maturity date. Assume that 10 policies are randomly selected from the policies data base


What is the probability that:


No more than 3 of these 10 insurance policies will have been surrendered before their maturity date? No more than 3 orders will be recieved?

1
Expert's answer
2022-03-01T07:16:47-0500

Let "X=" the numbers of insurance policies surrendered (cashed in) before their maturity date: "X\sim Bin (n, p)."

Given "n=10, p=0.2,q=0.8"


"P(X\le3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=\dbinom{10}{0}(0.2)^0(0.8)^{10-0}+\dbinom{10}{1}(0.2)^1(0.8)^{10-1}"

"+\dbinom{10}{2}(0.2)^2(0.8)^{10-2}+\dbinom{10}{3}(0.2)^3(0.8)^{10-3}"

"=0.1073741824+0.268435456"

"+0.301989888+0.201326592"

"=0.8791261184"